A hot plate 1 m × 1.5 m is maintained at 300°C. Air at 25°C blows over the

plate. If the convective heat transfer coefficient is 20 W/m2°C, calculate the rate of heat transfer


1
Expert's answer
2021-09-02T10:41:49-0400

Rate of heat transfer:

"Q = hAΔT \\

h=20 \;W/m^2 \cdot C \\

A= 1 \times 1.5 = 1.5 \;m^2 \\

ΔT = 300-25=275 \; °C \\

Q = 20 \times 1.5 \times 275 = 8250 \;W"

Answer: 8250 W